Commissioning Notes for 20080419 UT

Observer: D. Miller

Telescope Operator: A. Ceranski

Summary Of Evening

After a bit of collimation practice, the AGw unit was powered and the Guide Probe move to its nominal on-axis position. Star light was seen! The nominal on-axis position and focus position were fine. Around 9:45 the hexapod started acting up, and then buffer and RPC errors and then communication problems etc cause a couple of full TCS shutdowns and restarts. After about 2 hours the telescope was up and running again and we were back to the AGW.

The next step was to get light onto the WFSC. Again, the nominal position of the the AGw stage was fine (ie the hotspot is very close to where it was suppose to be). Several sets of WFSC images and ef pairs for the Guider camera and the IRTC were taken while seeing conditions were stable (~1"). This will allow three determinations of approximately the same wavefront so we can compare the EF curvature sensing results to the S-H WFSC.

For the last couple hours, collimation data was gathered to build a collimation table.

Overall, a very successful night.

DLM

Observing Log

NOTE:
  • Guider and WFSC data in directory /Repository/AGW_Data/080419
  • IRTC data (single: irtc2_* cubes: ef_*) in directory /Repository/IRTC_Data/080419

Collimation Star M5_0354: Elevation ~47

Added -2000 Z11 and 2000 Z6 to primary. Image on IRTC looks great (4pixelsFWHM??)
  • image 033543, 034231, 034316 focus change of -0.05 on primary 034612
  • integration time 0.0001 sec

0.5 focus on secondary integration time 0.001

08:57 MST inex_seq.cl /Repository/EF/080419/jose_00027_cube.fit 00028 00029

Started up AGw server. UMAC was cold ~7C and would not move the motors. About 6-7 minutes later its temperature was up to ~10.5 and I could home/move the motors. This will be a problem in the winter

Move Guide probe to (0, 600)

9:31 MST Starlight on the AGw guide probe!!

==> This next apparent problem was not real, hexapod problem probably caused this

/Repository/AGW_Data/080419/Guider00002.fits. Way out of focus. Focus stage at 0.0. Nominal focus at 21.0

Focus at 21.0, still way out of focus 500+ pixel pupil

Focus at 47.0 (end of range) still 230 pixel pupil. Looks like IRTC and AGw Guider are not co-focal

Moved Hexapod to z = -0.8 but still not focused. The move from -0.3 to -0.8 seems to have shifted the star image over about 200 pixels 0010 to 0011 fits images.


21:44

Requested move to z = -0.7 and Hexapod got into some strange open loop state and would try to move and then stop. Aaron called Paul, whose comment was "Hmmm, that's strange". He suggested that Aaron perform an "Init" on the Hexpod. This should be done anytime the hexapod looses power.

More error from OSS: RPCERROR and invalid buffer.

23:45 Telescope is now working

Guider

  • Pointed at a dimmer star "WT10_226"

  • Saved vibration data /Repository/EF/080419/vibration_00030_cube.fit (See Issue #1284)

  • Out of focus image z = 0.5 guider000032.fits: coma

  • Out of focus image z = 0.5 -800 z8 to remove coma guider000035.fits
    • Good focus guider000036.fits
    • z = -0.5 guider000037.fits

  • Moved star to x~292 y~255
    • First WFSC image Looks great, not much aberration wfsc000002.fits
    • guider000047.fits (guider a bit out of focus)

  • Good focus z = -0.12 on secondary guider000058.fits, wfsc000003.fits wfsc000004.fits guider0000059.fits

  • Good focus z = 0.07 on primary z=0 on secondary

  • EF pair z=0.5 on secondary guider0000064.fits (with coma) z=-0.5 on secondary guider000065.fits

  • Good focus guider 000074.fits, wfsc000007.fits, 8, 9, 10

  • Good focus guider000080 wfsc000011, 12

  • guider000084 wfsc000013, 14, 15

  • EF pair z=0.5 guider000089, 90 z = -0.5 guider000091, 92

IRTC

  • 02:11 MST - EF pair ef_00031.fit 32, 33

Guider

  • guider000096.fits wfsc000016, 17, 18 guider000097

  • ef pair z = 0.5 guider000098, 99 z = -0.5 guider000100, 101

IRTC

  • 02:23 MST ef pair ef_000034_cube.fit, 35, 36

M92

  • too dim to collimate, maybe 9th mag stars guider000102.fits, 107

M13

  • some collimation z11, z8, z6

  • good image guider000125.fit, 126 127
    • sky (not void of stars) 128, 129, 130(best)

  • irtc2...104300, 104527 sky 104756 [JMH: very few unsaturated stars, but images appear to be 0.4-0.5 arcsec FWHM.]

Collimation Data

All files are of the form irtc2_20080419.Time.fits in directory /Repository/IRTC_Data/080419

Elevation UT Time
79 112553
74 113356
60 114043
50 114547
34 115026
27 115720
34 120312
48 121129
55 121921
65 122347

-- DougMiller - 19 Apr 2008
Topic revision: r10 - 20 Apr 2008, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback